Home‎ > ‎Crew 143‎ > ‎

1117-EngineeringReport

Engineering Report
11/17/2014
Ian Siversides

A defogging device consisting of magnets placed in socks on each side of a helmet was tested during today’s EVA with mitigated success. The device worked until one of the magnets shifted in the sock and was no longer controlled by the other magnet. This idea has been improved upon using cloth tightly sewn around the magnets. Also, having a sock inside the helmet became a choking hazard.

The pool of water on the south side of the Hab reported in 1116 engineering report was filled with regolith. Rainwater is believed to be the source.

All ATVs started except for ATV 3 blue. DG was heard testing the ATVs mid afternoon.

The headset cable which previously broke (see 1114 engineering report) was damaged again during EVA.
This headset has been repaired.

During EVA, measurements were taken for an ATV shelter and for completing the tunnel to the observatory. Awaiting instructions for a construction project.

The air vents from the crew quarters merge into a main vent ending near the ceiling central air hole. This vent has a clappet valve which is in the closed position, blocking air circulation in the rooms when the doors are closed. The lack of air circulation results in hot and stuffy rooms in the morning. A toiler paper roll was used to prop the clappet valve open.

Trailer water tank: 30% (all other tanks full)
Propane tank: 23%, propane smell noticed through helmet. As discussed with Chuck, this can be expected.
Diesel tank: 75%